$SPAD/src/input file.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool file.output
)set message test on
)set message auto off
)clear all
--S 1 of 12
ifile:File List Integer:=open("/tmp/jazz1","output")
--R
--R
--R   (1)  "/tmp/jazz1"
--R                                                    Type: File(List(Integer))
--E 1
--S 2 of 12
write!(ifile, [-1,2,3])
--R
--R
--R   (2)  [- 1,2,3]
--R                                                          Type: List(Integer)
--E 2
--S 3 of 12
write!(ifile, [10,-10,0,111])
--R
--R
--R   (3)  [10,- 10,0,111]
--R                                                          Type: List(Integer)
--E 3
--S 4 of 12
write!(ifile, [7])
--R
--R
--R   (4)  [7]
--R                                                          Type: List(Integer)
--E 4
--S 5 of 12
reopen!(ifile, "input")

--R
--R
--R   (5)  "/tmp/jazz1"
--R                                                    Type: File(List(Integer))
--E 5
--S 6 of 12
read! ifile
--R
--R
--R   (6)  [- 1,2,3]
--R                                                          Type: List(Integer)
--E 6
--S 7 of 12
read! ifile
--R
--R
--R   (7)  [10,- 10,0,111]
--R                                                          Type: List(Integer)
--E 7
--S 8 of 12
readIfCan! ifile
--R
--R
--R   (8)  [7]
--R                                               Type: Union(List(Integer),...)
--E 8
--S 9 of 12
readIfCan! ifile
--R
--R
--R   (9)  "failed"
--R                                                    Type: Union("failed",...)
--E 9
--S 10 of 12
iomode ifile
--R
--R
--R   (10)  "input"
--R                                                                 Type: String
--E 10
--S 11 of 12
name ifile
--R
--R

--R   (11)  "/tmp/jazz1"
--R                                                               Type: FileName
--E 11
--S 12 of 12
close! ifile
--R
--R
--R   (12)  "/tmp/jazz1"
--R                                                    Type: File(List(Integer))
--E 12
)spool
)system rm /tmp/jazz1
)lisp (bye)

———-

References^{[1] nothing}

6